OpenStudy (anonymous):

If x = 2+square root of 3 divided by 2-square root of 3, find: x-1 over x

OpenStudy (whpalmer4):

For messy problems like this, I like to substitute \[a = \sqrt{3}\]do the algebra, then substitute back at the end, after the dust has settled. Your mileage may vary.

OpenStudy (tanner23456):

\[x=(2+\sqrt{3})/(2-\sqrt{3})\] multiply conjugate -> 2+sqrt(3) to top and bottom \[=(2+4\sqrt{3}+3)\] plugin for x-1/x \[((2+4\sqrt{3}+3)-1)/(2+4\sqrt{3}+3) =( 4+4\sqrt{3})/(5+4\sqrt{3})\] multiply conjugate again -> 5-4sqrt(3) to top and bottom = \[(4\sqrt{3}-28)/53\] = aprox. -0.398

OpenStudy (anonymous):

the answer is 8 multiply square root of 3

OpenStudy (tanner23456):

Well, can't shoot a guy for trying. It's been a while

OpenStudy (anonymous):

thanks for trying

OpenStudy (anonymous):

simplifying, you should get \(\large x=7+4\sqrt3 \)

OpenStudy (anonymous):

now plug that in for (x-1)/x

OpenStudy (anonymous):

and simplify again....

OpenStudy (agent0smith):

\[x=\frac{ (2+\sqrt{3})}{(2-\sqrt{3}) } \times \frac{ (2+\sqrt{3}) }{(2+\sqrt{3}) } = \] You'll get what @dpaInc posted.

OpenStudy (anonymous):

right i know that part but then i was stuck on the second part. i didnt know i had to plug it

OpenStudy (agent0smith):

\[\frac{ x -1 }{ x } =\frac{ 7+4\sqrt3-1 }{ 7+4\sqrt3 } \times \frac{ 7-4\sqrt3 }{ 7-4\sqrt3 }\]

OpenStudy (agent0smith):

^ again, I multiplied by the conjugate of the denominator \[\frac{ 6+4\sqrt3 }{ 7+4\sqrt3 } \times \frac{ 7-4\sqrt3 }{ 7-4\sqrt3 } = \frac{ 42+28\sqrt3-24\sqrt3-16*3 }{49-16*3 }\]

OpenStudy (anonymous):

i have the answer as 8 multiplied by square root of 3

OpenStudy (agent0smith):

hmm... maybe I've made a mistake.

OpenStudy (agent0smith):

Nope, my answer is correct.

OpenStudy (anonymous):

i have the answer in my textbook as 8 multiplied by square root of 3

OpenStudy (agent0smith):

Are you sure you wrote the problem correctly? is it \[x=\frac{ (2+\sqrt{3})}{(2-\sqrt{3}) } \]

OpenStudy (agent0smith):

Check it on a calculator. x = 13.928 Then find\[\frac{ x-1 }{ x }\] and you'll get 0.928. Then check \[8\sqrt3 = \] and you'll get 13.856.

OpenStudy (anonymous):

ill check it just give me a sec

OpenStudy (anonymous):

they are almost the same. im wondering where 8 multiplied by square root of 3 came from

OpenStudy (agent0smith):

They aren't almost the same. x is close to 8sqrt3... but (x-1)/x is not even close to 8sqrt3.

OpenStudy (anonymous):

i believe this problem can only be solved by a calculator as it is a digital worksheet. sorry made a mistake